- Remote participants by eZuce: Miyoki
Report of work (*:responsible person)
- Alignment, ETMY: *Enomoto, Miyo, Miyamoto
- Tuning of ETMY damping. Yaw of 0.2Hz of ETMY is still moving. RMS is limited by this ~10urad.
- ETEY had jump during weekend.
- BS has also jump on alignment, but not related duct connection.
- Duct connection made some jumps
- PR3 had jump 11am on Sunday.
- IMMT oplev was not centered before duct connection.
- CRY: Mirror reached at 18K.
Work plan (*:responsible person)
- VAC: Uchiyama
- Duct connection
- phi800 door of IX will be closed today.
- phi800 door of IY will be closed tomorrow or day after tomorrow.
- [AM] IR: PSL, Green: off.
- [PM] IR: PSL, Green: off.
- week of 3/12
- 3/14: meeting for pump-down check
- TMS installation at ETMY.
- Connecting ducts at the corner area.
- week of 3/19
- Vacuum test at at the corner area.
- TMS alignment using green light at ETMY.
- Beam alignment to REFL port.
- Connection of PD signal to DAQ should be implemented this week or next week.
- GigE camera should be placed at the REFL port eventually.